OpenStudy (anonymous):

MEDAL AND FAN!!! Gem counted 54 wheels on 3 semitrucks. The same number of wheels were on each semitruck. How many wheels were on each semitruck? A.54 over 3 B. 54 • 3 C. 54 – 3 D. 3 + 54 @darkb0nebeauty

OpenStudy (anonymous):



OpenStudy (anonymous):

@danish071996 @iceicebaby

OpenStudy (anonymous):

@darkb0nebeauty

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

Hang on wrong one

OpenStudy (anonymous):

i say A!! not 100% but im 99.98% sure

OpenStudy (anonymous):

The Mendaz family's car traveled 196 miles for every 7 gallons of gasoline that it used. Which equation can be used to determine how many miles their car traveled on each gallon of gasoline?

OpenStudy (anonymous):

uh okay

OpenStudy (anonymous):

fan and medal

OpenStudy (anonymous):

Hang on and thank you

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and no problem

OpenStudy (anonymous):

A grocery store is selling 12 bannanas for $2.40. What is the best cost for 15 bannanas? a.)$2.52 b.)$2.55 c.)$3.00 d,)$36.00 will fan and medal

OpenStudy (anonymous):

ASAP

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

@darkb0nebeauty

OpenStudy (anonymous):

@xo.minnie.xox

sammixboo (sammixboo):

Find the cost for 1 banana by dividing 12/2.4

OpenStudy (anonymous):

C is the answer

OpenStudy (anonymous):

@IndiiAvery

OpenStudy (mathstudent55):

54 wheels on 3 semi trucks means 54/3 = 18 wheels in each truck. 196 miles using 7 gallons means 196/7 = 28 miles per gallon. 12 bananas for $2.40 means $2.40/12 = $0.2 per banana. $0.2 * 15 = $3.00 for 15 bananas